A box rests on a frozen pond, which serves as a frictionless

horizontal surface. If a fisherman applies a horizontal force with magnitude 48.0 N to the box and produces an acceleration of magnitude 3.00 m/s², what is the mass of the box?


1
Expert's answer
2021-12-12T16:42:40-0500

If the fisherman applies a horizontal force of magnitude 48 N , the produced acceleration on the box measures to be = 3 m/s2

thus from Newton's second law we write

Net force "= mass \times acceleration"

"F = m \times a"

mass of the box "= \frac{Force }{ acceleration}"

"= \frac{48 }{ 3}"

= 16 kg